Can someone solve question 6 for me? Thank you very much.

Part 1: Rigged Bids as a Price Ceiling Suppose there are five shale energy companies bidding on leases to drill for oil on government land. A lease gives a company the right to drill for oil in the designated land. Since each company has a different cost of drilling, each company receives a different level of benefit from getting a lease. The benefits are as follow: Company Profit from Lease (Millions of $) Chevron Encana Hess Exxon Mobile 7 10 Meanwhile, there are five state governments with shale fields to lease. Each state incurs a different cost (in environmental damage, for example) for leasing out its field. The costs are as follow: State Cost from Leasing (Millions of S) Texas New Yorlk Pennsylvania Oklahoma 9 Who is the supply side in the market for leases? Who is the demand side? Reorder the tables above so that the demanders are in descending order of willingness-to-pay, and the suppliers are in increasing order of cost. What is the market-clearing price of a lease? How many leases are sold? Draw the supply and demand curves in the market for leases. Label the market-clearing price and quantity What is the consumer surplus? What is the producer surplus? What is total surplus? 1. 2. 3. 4. Now suppose the oil firms agree to a scheme in which they do not bid more than $4 million. This acts like a price ceiling of $4 million per lease 5 Now what is the price of a lease? How many leases are sold? Is there excess demand? How many demanders do not get a lease? Now what is consumer surplus? What is producer surplus? Who has benefited and who has lost? Has total surplus risen, fallen, or stayed the same? Is this what we should expect in general as the effect of a price ceiling? Draw the price ceiling on your graph of the supply and demand curves 6. 7. 8. 1 If you have excess demand, assume the consumers with the lowest willingness-to-pay do not get a lease. If you have excess supply, assume the sellers with the highest cost do not sell a lease.
Answer #1

Producer surplus (PS)is defined as the difference between how much a producer is willing to sell his goods for and the amount he receives from selling the good.

Consumer surplus(CS) is defined as the difference between how much a consumer is willing to pay for a good and how much he is actually paying.

Price ceiling is the price above which price of a good cannot go. It is upper bound. For question 6,

the following states and companies have the following PS and CS

COMPANIES CS($ mil) STATES PS ($ mil)

Chevron 7-4=3 Texas 4-3=1

Encana 5-4=1 NY 4-9=-5

Hess 3-4=-1 PENN 4-5=-1

Exxon 10-4=6 OKLAHOMA 4-1=3

So we can see that consumers(companies) Chevron, encana and exxon have benefitted as their profit generated will be more than what they are having to pay to lease the land. Exxon benefits the most(6 million dollars). Hess is the only company that losses as it will suffer a loss of 1 million dollars because it can only profit 3 million dollars which is not enough to recover the cost.

In case of producers(states), Texas and oklahoma benefits as they will receive more money than the cost they incur. Okhlahoma benefits most(3 million dollars). But NY and Pennsylvania incurs more cost than 4 million dollars and thus will suffer loss with NY suffering the most(5 million dollars).